OpenStudy (anonymous):

Sovle 3(x-4)-5x= 2x=9

OpenStudy (anonymous):

first distribute the 3 and what do u get?

OpenStudy (anonymous):

3x-12

OpenStudy (anonymous):

yes now you have 3x-12-5x = 2x is that supposed to be a + or - where the equals is after 2x ?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

ok :) so you have 3x-12-5x+2x+9 correct now for the first half 3x-12-5x <---- combine like terms so 3x-5x=?

OpenStudy (anonymous):

oops 3x-12-5x=2x+9 lol

OpenStudy (anonymous):

8x-12=2x+9

OpenStudy (anonymous):

oh wait -2x-5

OpenStudy (anonymous):

yes -2x-5=2x+9 now if you subtract the 2x from the -2x what happens then

OpenStudy (anonymous):